# Unifying Search and Recommendation: A Generative Paradigm Inspired by Information Theory

### v2

@@ 4.1.2. S&R history soft filtering
where $\alpha_{j}$ is the importance score for the $j$-th historical item, and $\text{MCA}(e_{i_{c}},h_{j})$ represents the attention score between the candidate item’s CF embedding $e_{i_{c}}$ and the $j$-th item CF embedding $h_{j}$.
@@ 4.2.1. Training
where $\beta$ is the loss balance factor that controls the weight of the contrastive loss relative to the S&R task loss, and $|y|$ is the length of the target sequence, $y_{t}$ is the token at position $t$ in the target sequence, and $y_{<t}$ represents all tokens preceding $t$. The conditional probability $P_{\Theta}(y_{t}\mid y_{<t},E)$ is computed by the generative model, parameterized by overall model parameters $\Theta$, given the input tokenized embedding $E$. This term encourages the model to maximize the likelihood of generating the correct output sequence for S&R.
@@ 4.2.2. Contrastive learning between CF and Semantic views
5. Mutual Information Analysis of GenSR
After illustrating the detailed methodology of GenSR, we conduct an in-depth analysis of mutual information to support our claim that the proposed generative paradigm effectively enhances mutual information compared to traditional discriminative methods. We begin by defining mutual information from an information-theoretic perspective, and then present both theoretical and empirical evidence to substantiate this claim in the following sections.
5.1. Information-theoretic perspective
From an information-theoretic standpoint, the representational capability of a unified S&R model can be examined via the mutual information between the input $X$ and the output $Y_{t}$ for a task $t\in\{S,R\}$. Mutual information, denoted as $I(X;Y_{t})$, measures how much information the input $X$ provides about the output $Y_{t}$, quantifying their dependence. A higher mutual information indicates that the input $X$ contains more predictive information about the output $Y_{t}$, reflecting a stronger alignment between inputs and task-specific outputs, i.e., higher model representation capability. Conversely, lower mutual information implies that predicted output fail to fully capture the distribution and pattern of input, reflecting a lower model representation capability. Therefore, mutual information $I(X;Y_{t})$ serves as a good indicator for model representation capability (Bachman et al., 2019).
5.2. Theoretical analysis
In this section, we aim to show the effectiveness of our model GenSR, based on the aforementioned mutual information. Specifically, we conduct a theoretical analysis to show this utilizing the following lemma:
Lemma 5.1.
Let the input $X\sim\mathcal{N}(\mu,\Sigma)$, and assume that noise $\epsilon\sim\mathcal{N}(0,\sigma_{\epsilon}^{2}I)$ is independent Gaussian noise. Consider two paradigms: generative (with parameters $\phi$) and discriminative (with parameter $\theta$). The mutual information between the input and output satisfies:
We provide a detailed proof of Lemma 5.1 from the view of parameter approximation separability in the Appendix. Intuitively, we prove that generative paradigm can enhance mutual information in S&R scenario compared with discriminative paradigm under certain input distribution assumption.
5.3. Empirical analysis
@@
To further support our theoretical claim and verify the conclusion in a broader complex real-world scenario, we employ Mutual Information Neural Estimation (MINE) to empirically estimate the practical mutual information between the input $X$ and the task-specific outputs $Y_{S}$ and $Y_{R}$ (Belghazi et al., 2018). Our empirical findings are consistent with the theoretical conclusion. Specifically, we use the Donsker-Varadhan representation of KL divergence to estimate the lower bound of mutual information as:

### v3

@@ 5.1.2. S&R history soft filtering
where $\alpha_{j}$ is the importance score for the $j$-th historical item, and $\text{MCA}(e_{i_{c}},h_{j})$ represents the attention score between the candidate item’s CF embedding $e_{i_{c}}$ and the $j$-th item CF embedding $h_{j}$. Through the MCA module, the candidate item $e_{i_{c}}$ dynamically queries and weights the most relevant historical item embeddings $h_{j}$, providing importance signals for subsequent soft filtering.
@@ 5.2.2. Contrastive learning between CF and Semantic views
5.2.3. Loss function
where $\beta$ is the loss balance factor that controls the weight of the contrastive loss relative to the S&R task loss, and $|y|$ is the length of the target sequence, $y_{t}$ is the token at position $t$ in the target sequence, and $y_{<t}$ represents all tokens preceding $t$. The conditional probability $P_{\Theta}(y_{t}\mid y_{<t},E)$ is computed by the generative model, parameterized by overall model parameters $\Theta$, given the input tokenized embedding $E$. Intuitively, generation loss $\mathcal{L}_{\text{S\&R}}$ encourages the model to maximize the likelihood of generating the correct output sequence for S&R.
6. Experiments
In this section, we conduct a comprehensive experimental study to address the following research questions.
  1. -
    RQ1: How does the performance of GenSR compare to other baseline methods across various datasets in both recommendation and search scenarios?
  2. -
    RQ2: How well can GenSR enhance mutual information compared to discriminative approaches?
  3. -
    RQ3: How effectively does GenSR mitigate gradient conflict relative to discriminative approaches?
  4. -
    RQ4: What is the impact of different strategies (e.g., dual-view representation and S&R history soft filtering) within GenSR on its overall performance?
Table 1. Statistics of three datasets. #Inter-S and #Inter-R denote the interaction number of search behavior and recommendation behavior, respectively.
Dataset #User #Item #Queries #Inter-S #Inter-R
Amazon 68,223 61,934 4,298 934,664 989,618
KuaiSAR 25,877 6,890,707 453,667 5,059,169 14,605,716
Table 2. Overall performance of GenSR and other baselines under recommendation scenario. Bold signifies the best performance among the baselines and GenSR while underline represents the second-best method. * denotes statistically significant improvements of GenSR over the best baseline, according to t-tests with a significance level of $p$ ¡ 0.01.
Amazon KuaiSAR
Method R@5 R@10 N@5 N@10 R@5 R@10 N@5 N@10
DIN (Zhou et al., 2018) 0.5170 0.6525 0.3726 0.4165 0.4509 0.6179 0.3104 0.3643
GRU4Rec (Hidasi et al., 2016) 0.4949 0.6548 0.3388 0.3907 0.3764 0.5788 0.2345 0.3087
SASRec (Kang and McAuley, 2018) 0.5295 0.6772 0.3747 0.4225 0.4065 0.6007 0.2671 0.3298
BERT4Rec (Sun et al., 2019) 0.5311 0.6658 0.3954 0.4390 0.3699 0.5885 0.2381 0.3083
FMLP-Rec (Zhou et al., 2022) 0.5356 0.6879 0.3739 0.4232 0.4292 0.6159 0.2851 0.3453
CoLLM (Zhang et al., 2023) 0.5800 0.6811 0.4651 0.5001 0.4092 0.5915 0.3199 0.3683
Recommendation IDGenRec (Tan et al., 2024) 0.2305 0.2818 0.1791 0.1955 - - - -
NRHUB (Wu et al., 2019) 0.4988 0.6503 0.3487 0.3977 0.3862 0.5610 0.2572 0.3136
Query-SeqRec (He et al., 2022) 0.5401 0.6860 0.3859 0.4333 0.3920 0.5890 0.2552 0.3186
SESRec (Si et al., 2023) 0.5623 0.6864 0.4245 0.4648 0.4956 0.6643 0.3432 0.3978
JSR (Zamani and Croft, 2018) 0.5467 0.6779 0.3970 0.4396 0.4791 0.6453 0.3315 0.3853
USER (Yao et al., 2021) 0.5441 0.6854 0.3964 0.4422 0.4086 0.5627 0.2820 0.3318
UnifiedSSR (Xie et al., 2024) 0.5196 0.6707 0.3662 0.4151 0.3981 0.5939 0.2617 0.3249
UniSAR (Shi et al., 2024) 0.5874 0.7020 0.4513 0.4885 0.5169 0.6792 0.3632 0.4158
BSR (Penha et al., 2024) 0.1417 0.2278 0.0899 0.1177 0.2180 0.3395 0.1472 0.1860
Joint S&R GenSR 0.6102* 0.7200* 0.4922* 0.5278* 0.5235* 0.6819* 0.3686* 0.4199*
Table 3. Overall performance of GenSR and other baselines under search scenario. Bold signifies the best performance among the baselines and GenSR while underline represents the second-best method. * denotes statistically significant improvements of GenSR over the best baseline, according to t-tests with a significance level of $p$ ¡ 0.01.
Amazon KuaiSAR
Method R@5 R@10 N@5 N@10 R@5 R@10 N@5 N@10
QEM (Ai et al., 2019) 0.7100 0.8186 0.5066 0.5422 0.6020 0.7182 0.4575 0.4953
HEM (Ai et al., 2017) 0.6778 0.8267 0.4736 0.5221 0.6505 0.7653 0.5029 0.5400
AEM (Ai et al., 2019) 0.7095 0.8443 0.5114 0.5554 0.5956 0.7182 0.4415 0.4812
ZAM (Ai et al., 2019) 0.7109 0.8468 0.5147 0.5590 0.6117 0.7344 0.4560 0.4959
TEM (Bi et al., 2020) 0.8185 0.9051 0.6303 0.6587 0.6502 0.7632 0.4887 0.5254
Search CoPPS (Dai et al., 2023) 0.8169 0.9051 0.6281 0.6570 0.6616 0.7707 0.4977 0.5331
JSR (Zamani and Croft, 2018) 0.7038 0.8225 0.5173 0.5563 0.7162 0.7961 0.5962 0.6221
USER (Yao et al., 2021) 0.7631 0.8697 0.6000 0.6348 0.7304 0.8149 0.6069 0.6342
UnifiedSSR (Xie et al., 2024) 0.7744 0.8812 0.5847 0.6196 0.7377 0.8320 0.5991 0.6297
UniSAR (Shi et al., 2024) 0.8190 0.8977 0.6875 0.7132 0.7476 0.8369 0.6417 0.6708
BSR (Penha et al., 2024) 0.4786 0.6015 0.3485 0.3883 0.1573 0.2075 0.1359 0.1520
Joint S&R GenSR 0.9782* 0.9853* 0.9664* 0.9687* 0.7661* 0.8372 0.7564* 0.7655*
6.1. Experimental setups
6.1.1. Datasets.
We evaluate GenSR on the two publicly available datasets that contain both S&R histories.
  1. (1)
    Amazon [^4] https://jmcauley.ucsd.edu/data/amazon/. is a publicly released product review dataset collected from Amazon. It contains extensive user-generated product reviews and metadata, such as product titles and descriptions. We use the ”Kindle Store” 5-core sub-dataset from the dataset, which contains user reviews and ratings for Kindle books. Adopting the procedure from prior work (Si et al., 2023; Shi et al., 2024), we first generate synthetic search behaviors and queries based on this recommendation dataset and then apply a leave-one-out strategy for dataset splitting. The detailed process for generating search behaviors is described in (Ai et al., 2017). Specifically, queries are extracted from the hierarchical category information of each item and paired with user purchase behaviors to construct user–query–item triples. Due to the limited availability of real-world unified search and recommendation datasets, this automatically generated search dataset has been widely adopted in prior research on unifying search and recommendation (Si et al., 2023; Shi et al., 2024; Ai et al., 2017; Van Gysel et al., 2016; Guo et al., 2019).
  2. (2)
    KuaiSAR [^5] https://kuaisar.github.io/. is a micro-video dataset that offers authentic S&R behaviors of users, featuring detailed item categories and real user queries. We preprocess the data to include only users showing both S&R behaviors. Following (Shi et al., 2024), we filter for users and items with at least five interactions and allocate the latest day’s data for testing, the previous day for validation, and the earlier data for training.
The statistics of datasets are shown in Table 1.
6.1.2. Baselines.
The recommendation baselines are:
  • DIN (Zhou et al., 2018) employs a local activation unit that adaptively learns user interest representations, weighting historical behaviors based on their relevance to a candidate ad.
  • GRU4Rec (Hidasi et al., 2016) utilizes RNNs for session-based recommendations. It models sequential user interactions within short sessions, improving recommendations where user profiles are absent.
  • SASRec (Kang and McAuley, 2018) introduces a self-attention based model for sequential recommendation, adeptly capturing long-term user behavior by adaptively focusing on relevant past items.
  • BERT4Rec (Sun et al., 2019) employs deep bidirectional self-attention with a Cloze objective to learn richer item representations by leveraging both left and right contexts in user behavior sequences.
  • FMLP (Zhou et al., 2022) uses an all-MLP model that employs learnable filters in the frequency domain to effectively denoise sequential user behavior data.
  • CoLLM (Zhang et al., 2023) integrates collaborative embeddings derived from external traditional recommender models into LLMs by mapping them to the LLM’s input token space.
  • IDGenRec (Tan et al., 2024) proposes learning unique, semantically rich textual IDs for items to seamlessly integrate them into LLM-based generative recommender systems. [^6] We only implement IDGenRec on Amazon due to its requirement for explicit semantic textual information of items, which is unavailable for KuaiSAR.
The search baselines are:
  • QEM (Ai et al., 2017) learns a dense vector representation for a given search query, allowing for semantic similarity calculations with products in the same embedding space.
  • HEM (Ai et al., 2017) jointly learns distributed representations for words, queries, users, and products in a unified latent space, effectively personalizing product search by composing user and query embeddings.
  • AEM (Ai et al., 2019) is an attention-based model that personalizes product search by dynamically weighting items in a user’s purchase history according to the current query.
  • ZAM (Ai et al., 2019) introduces a novel Zero Attention Strategy, allowing the model to automatically determine when and how much to personalize by optionally attending to a zero vector, thus enabling differentiated personalization.
  • TEM (Bi et al., 2020) leverages a transformer architecture to dynamically control the influence of personalization and model the interactions between purchased items for personalized product search.
  • CoPPS (Dai et al., 2023) introduces contrastive learning to learn robust user representations for personalized product search by leveraging diverse data augmentation strategies that exploit intrinsic sequence correlations and external knowledge graphs to overcome data sparsity and noise.
The unifying baselines include two categories:
(a) Search enhanced recommendation:
  • NRHUB (Wu et al., 2019) enriches user representations by integrating heterogeneous user behaviors (search queries, browsed webpages, and clicked news) through an attentive multi-view learning framework.
  • Query-SeqRec (He et al., 2022) leverages previously ignored user queries as a crucial contextual signal to enhance sequential recommendation accuracy by better understanding evolving user intent
  • SESRec (Si et al., 2023) enhances sequential recommendation by learning disentangled representations of similar and dissimilar user interests from their combined search and recommendation behaviors.
(b) Jointly unified S&R:
  • JSR (Zamani and Croft, 2018) jointly models search and recommendation tasks by learning shared item representations and minimizing a combined loss function.
  • USER (Yao et al., 2021) integrates heterogeneous user search and recommendation behaviors into a single chronological sequence to learn comprehensive user interests and leverage the interdependencies between the two tasks.
  • UnifiedSSR (Xie et al., 2024) leverags cross-scenario and cross-view information sharing and dynamically modeling user intent through self-supervised session discovery.
  • UniSAR (Shi et al., 2024) introduces a novel framework to explicitly extract, align, and fuse various fine-grained user transition behaviors between search and recommendation.
  • BSR (Penha et al., 2024) employs atomic item IDs to jointly train generative models for search and recommendation.
6.1.3. Experimental settings.
Following previous work (Shi et al., 2024; Xie et al., 2024; Yao et al., 2021), we randomly sample 99 negative items that the user has not interacted with and combine them with the ground truth item to create the candidate item list. All the models are then tasked with ranking these candidate items to evaluate top-K S&R performance using Recall@K and NDCG@K with $K=\{5,10\}$. For all the baselines, the hyper-parameters are searched according to the reporting range in the original papers. To ensure a fair comparison, we implement constrained beam search during inference for two full-ranking generative baselines, IDGenRec and GenSR, which guarantees that the generated results are restricted to the valid candidate set. Regarding our proposed GenSR, we select flan-t5-base as the backbone. The best hyper-parameters are selected with the searching scopes as follows: the temperature factor $\tau$ and the loss balance factor $\beta$ are tuned with the ranges $\{0.01,0.03,0.05,0.07,0.09\}$ and $\{0.001,0.01,0.1,0.15,0.2,0.25,0.3\}$.
6.2. Overall performance (RQ1)
We conducted a comprehensive experimental comparison of GenSR’s performance against various baseline methods in both recommendation and search scenarios. As shown in Table 2 and Table 3, the experimental results reveal several key findings:
  1. (1)
    Superiority of GenSR: GenSR consistently outperforms all baseline methods across all datasets in both S&R tasks, achieving state-of-the-art results. This good performance is primarily attributed to GenSR’s generative paradigm and its two core modules: Dual Representation Learning and S&R Task Unifying. As discussed in Section 4, the generative paradigm enhances the mutual information between input features and task outputs by constructing effective task-specific subspaces. This approach mitigates the gradient conflict issues commonly observed in traditional discriminative multi-task learning, as further illustrated in Sections 6.4.2 and Sections 6.5. After that, by preparing effective dual-view representations for each subspace and jointly optimizing them within one unified model, GenSR improves the performance of both tasks while fully leveraging the capabilities of a shared pre-trained model.
  2. (2)
    General Advantages of the Unified S&R Approach: We observe that unified S&R methods consistently outperform models designed for a single task, both recommendation-only models (e.g., DIN and SASRec) and search-only models (e.g., QEM and TEM). These findings provide strong evidence for the intrinsic value of integrating S&R within a shared framework. User behaviors in search and recommendation are not isolated but complementary, with each offering distinct and informative signals. Joint modeling enables a more comprehensive understanding of user preferences and facilitates richer item-side interaction signals, ultimately leading to mutual performance gains across both tasks.
  3. (3)
    Analysis of GenSR’s SOTA Performance on the Amazon Search Task: Notably, GenSR achieves particularly strong results on the Amazon dataset’s search task. This can be explained by the dataset’s nature: search queries in Amazon are synthesized from item descriptions, indicating a high degree of alignment and strong semantic correlation between queries and item content. Note that the use of synthetic queries for this dataset to use it for search purposes is in line with prior work (see Section 6.1.1). Thanks to GenSR’s dual representation learning module, particularly the semantic-view, GenSR is exceptionally capable of capturing and utilizing such fine-grained semantic matching. When queries are semantically aligned with item descriptions, GenSR’s generative capabilities allow it to more precisely infer user intent and retrieve relevant items, thereby yielding remarkable performance improvements. This highlights the generative paradigm’s strong potential in leveraging semantic information, especially on semantically rich and well-aligned datasets.
  4. (4)
    Limitations of Existing Baseline Methods: Not all generative or unified models can achieve GenSR’s performance level. For example, IDGenRec performs relatively poorly, especially when compared with GenSR. This is likely due to its heavy reliance on semantic textual information to generate unique item IDs, which may hinder its ability to fully capture and utilize traditional CF signals in recommendation tasks. In many cases, pure semantic matching is insufficient to reflect complex user-item collaborative preferences. GenSR addresses this issue through dual representation learning. Similarly, BSR exhibits significantly limited performance. By generating atomic item IDs, BSR lacks enriched semantic information to enhance item representations. Relying solely on atomic IDs may result in poor modeling of item properties and deep user preferences, especially when semantic matching is crucial for search tasks.
6.3. Potential Application to Full Ranking (RQ1)
Full ranking represents a highly challenging yet critical evaluation setting for real-world S&R systems. Unlike the more common candidate re-ranking evaluation scenario in unified S&R, where models operate on a small, pre-selected set of candidate items, full-ranking requires identifying the most relevant items directly from the entire corpus of available items. Success in full-ranking demonstrates a model’s robust capability for open-vocabulary item discovery and global relevance estimation, pushing the limits of what search engines and recommender systems can achieve.
A key advantage of the generative paradigm in S&R lies in its inherent ability to perform full-ranking, seamlessly unifying the retrieval and re-ranking stages into a single, end-to-end optimization framework (Sun et al., 2023). Unlike traditional two-stage pipelines which first retrieve a subset of candidates and subsequently re-rank them, a generative model directly produces the final ranked items in a single pass. This unified approach not only simplifies the overall architecture but also allows the model to jointly optimize retrieval and re-ranking without relying on intermediate heuristic filtering. To comprehensively assess the global ranking capability of this generative paradigm, we further evaluate GenSR in the full-ranking setting.
To adapt GenSR for the full-ranking task, we slightly modify its input and output strategy which is suitable for the more typical evaluation setting in unified S&R. In the standard evaluation setup, GenSR is typically given a candidate item $i_{c}$ along with user history and query, and it generates token probability to indicate preference. However, in the full-ranking scenario, the model’s goal shifts from assessing a pre-existing candidate to identifying the most relevant item from the entire corpus. To accommodate this, the input prompt to the generative model is adjusted by excluding the embeddings of the candidate item ($e_{i_{c}},s_{i_{c}}$). Instead, the model is provided only with the user’s interaction history $H_{u}$ and query $q$ (for search tasks), and it is trained to directly generate the relevant item ID through dual representation learning (excluding soft-filtering strategy due to the lack of candidate item) and S&R task unifying.
This adjustment transforms the task into an item ID generation problem, similar to current generative retrieval and recommendation methods. During inference, we employ constrained beam search, which ensures that the output consists of valid item IDs from the entire corpus. This approach effectively performs full-ranking, enabling the model to discover the most relevant items without relying on a pre-selected candidate set.
Crucially, this adaptation strategy aligns the computational complexity of GenSR for full-ranking with that of existing typical generative retrieval models (Wang et al., 2021; Penha et al., 2024). The primary computational burden lies in the decoding phase, where the model generates item IDs. During this process, constrained beam search is applied over the vocabulary of item IDs, with complexity approximately proportional to $B\times L\times V_{item}$. Here, $B$ is the beam size, $L$ is the length of the generated ID sequence (typically 1 for a single item ID), and $V_{item}$ is the total number of items in the corpus.
Table 4. Overall performance of GenSR and other baselines in the full-ranking setting. Bold signifies the best performance among the baselines and GenSR, and underline represents the second-best method. * denotes statistically significant improvements of GenSR over the best baseline, according to t-tests with a significance level of $p$ ¡ 0.01.
Amazon KuaiSAR
Method R@5 R@10 N@5 N@10 R@5 R@10 N@5 N@10
Recommendation TIGER (Rajput et al., 2023) 0.0108 0.0173 0.0101 0.0142 0.0149 0.0213 0.0138 0.0201
BSR (Penha et al., 2024) 0.0027 0.0037 0.0016 0.0019 0.0052 0.0081 0.0043 0.0071
GenSR 0.0153* 0.0211* 0.0136* 0.0198* 0.0187* 0.0254* 0.0181* 0.0250*
Search GenRet (Sun et al., 2023) 0.0221 0.0311 0.0197 0.0277 0.0112 0.0181 0.0099 0.0173
BSR (Penha et al., 2024) 0.0072 0.0080 0.0058 0.0059 0.0074 0.0109 0.0066 0.0095
GenSR 0.0268* 0.0362* 0.0223* 0.0305* 0.0181* 0.0259* 0.0144* 0.0216*
We compare GenSR with three representative generative models under the full-ranking setting:
  • BSR (Penha et al., 2024) is a generative joint S&R model that outputs items as plain IDs, without leveraging content semantics.
  • TIGER (Rajput et al., 2023) is a generative recommendation model that assigns each item a discrete Semantic ID and trains a sequence-to-sequence model to predict the next item’s Semantic ID in a user’s history
  • GenRet (Sun et al., 2023) is a generative retrieval model that learns to tokenize documents into short discrete identifiers and generate these identifiers for a given query.
These baselines include both unified S&R models and single-task generative approaches. We evaluate all models on the Amazon and KuaiSAR datasets under the full-ranking setting, using the same top-$K$ metrics as reported in Section 6.1. The results are presented in Table 4. As expected, all models experience a performance drop compared to the main setting due to the significantly larger candidate space. Nevertheless, GenSR consistently outperforms the generative baselines by a substantial margin across all metrics and datasets, owing to its superior dual representation learning and unified training strategy. Overall, these results strongly highlight GenSR’s remarkable generalization ability across different evaluation scenarios. While the typical candidate re-ranking setting remains widely used in unified S&R task, GenSR also demonstrates its capacity to perform effectively in the generative full-ranking setting. This underscores GenSR’s effectiveness and its potential to serve as a robust generative paradigm in search and recommendation tasks.
6.4. Mutual Information Analysis (RQ2)
In this section, we evaluate mutual information from two complementary angles. First, we provide quantitative evidence by estimating the mutual information gains directly. Second, we offer a qualitative geometric view via parameter-space visualization, showing how GenSR’s subspaces disentangle task manifolds and better align inputs with task-specific outputs, thereby increasing the mutual information between them.
6.4.1. Quantifying mutual information gains
@@
To further support our theoretical claim directly and verify the conclusion in a broader complex real-world scenario, we employ Mutual Information Neural Estimation (MINE) to empirically estimate the practical mutual information between the input $X$ and the task-specific outputs $Y_{S}$ and $Y_{R}$ (Belghazi et al., 2018). Our empirical findings are consistent with the theoretical conclusion. Specifically, we use the Donsker-Varadhan representation of KL divergence to estimate the lower bound of mutual information as: